OpenStudy (anonymous):

A= 2(L+W) Solve for W

OpenStudy (anonymous):

A=2L+2W 2W=A-2L W=(A-2L)/2

OpenStudy (anonymous):

Step by step please

OpenStudy (lin.ivory):

There are too variables in this equation, are you looking for a numerical answer?

OpenStudy (lin.ivory):

many*